# Radial Transport in High-Redshift Disk Galaxies Dominated by Inflowing Streams

### v1

@@ 3.2 General Trends
Having discussed a particular case, we now describe some general trends. In Fig. 3, the left- and right-hand panels, respectively, show the medians (solid lines) and $16^{\rm th}-84^{\rm th}$ percentiles (envelopes) of cold gas $V_{r}$ (in units of $V_{\rm rot}$) and $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$) as functions of $r$ (in units of $R_{\rm d}$) obtained from our disk galaxy sample. The median $V_{r}$ and $F_{r}$ are always negative and increase in magnitude with increasing $r$, except within $\sim\!0.2\ R_{\rm d}$, where the trends are reversed. While both $F_{r}$ and $V_{r}$ are positive at the $84^{\rm th}$ percentile for all $r$, their magnitudes at the $16^{\rm th}$ percentile are much larger. For example, at $r=R_{\rm d}$, $V_{r}$ at the $16^{\rm th}$ percentile is about $-0.25\ V_{\rm rot}$, while that at the $84^{\rm th}$ percentile is only $\sim\!0.04\ V_{\rm rot}$. Similarly, at the same radius, $F_{r}$ at the $16^{\rm th}$ percentile is about five times larger in magnitude than that at the $84^{\rm th}$ percentile. Therefore, in a statistical sense, the radial transport of cold gas in our disk galaxy sample is inflow-dominated at all radii, i.e., most of the mass is inflowing rather than outflowing.
In Fig. 4, the ratio of the disk-averaged radial velocity to the disk-averaged rotational velocity, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$, is plotted against $f$ and $z$ in the left- and right-hand panels, respectively, with the blue, round points. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $f$ and $z$ bins of varying widths (to include a sensible number of snapshots), as indicated by the horizontal error bars, with the locations of the squares along the $f$- and $z$-axes being at the bin-averages. We find that for most of the VELA disks ($82\%$), $\langle V_{r}\rangle$ is negative, again highlighting the inflow-dominated nature of cold gas radial transport in our disk galaxy sample (note that $\langle V_{\rm rot}\rangle$ is always positive). Also, on average, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ decreases (increases in magnitude) with an increase in both $f$ and $z$, indicating that galaxies with larger cold gas fractions or at higher redshifts have higher levels of disk-averaged radial inflow in general. Both trends are relatively weak, though the one with cold gas fraction is more pronounced than that with redshift, the Spearman’s rank correlation coefficients in the two cases being $-0.26$ and $-0.13$, respectively.
@@ 4.1 Viscous Disk Transport
Here, $\mathscr{L}$ is the rate at which turbulent kinetic energy per unit area is dissipated, $V$ is the circular velocity at radius, $r$, $\beta=d{\rm ln}V/d{\rm ln}r$, $\sigma$ is the $1$D gas velocity dispersion, assumed to be isotropic, and $\sigma_{\rm sf}$ is the value of $\sigma$ that can be maintained with star formation feedback alone without any additional energy input from radial transport of gas ($\dot{M}_{\rm in}=0$). Furthermore, $\dot{M}_{\rm in}=2\pi r\Sigma V_{\rm in}$, where $\Sigma$ and $V_{\rm in}$ are the gas surface density and inflow velocity, respectively. Therefore,
@@
where $\sqrt{\sigma^{2}-\sigma^{2}_{\rm th}}$ is the $1$D turbulent velocity dispersion of gas, $\sigma_{\rm th}$ being the $1$D thermal velocity dispersion, typically much smaller than $\sigma$, and $t_{\rm cross}=h/\sqrt{\sigma^{2}-\sigma^{2}_{\rm th}}$. The gas scale height is estimated following Forbes et al. (2012) as
@@
where $Q$ is the two-component Toomre-Q parameter, expressed as $Q=f_{g,Q}Q_{g}$, where $f_{g,Q}$ depends on the gas and stellar surface densities and velocity dispersions (see Equation 9 in K18, ), $\phi_{Q}=1+Q_{g}/Q_{*}$, and $\phi_{\rm nt}=1-\sigma_{\rm th}^{2}/\sigma^{2}$. Using Equations 9 and 11,
@@ 4.2 Ring Migration
where $\sigma$ is the average $1$D gas velocity dispersion, assumed to be isotropic, the angular velocity is approximated as
@@ 4.3 Clump Migration
Figure 5: Comparison of Disk Instability Based Radial Transport Models with Simulations: For each snapshot in our disk galaxy sample, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ is plotted against $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ with the blue, round points. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ bins of varying widths, as indicated by the horizontal error bars, with the locations of the squares along the $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ axis being at the bin-averages. The different curves show the inflow velocity predictions from the disk-instability-based models of radial transport, as labeled, with $Q=0.68$ (marginal instability condition for thick disks; Goldreich & Lynden-Bell, 1965) in the left-hand panel and $Q=1$ (marginal instability condition for thin disks) in the right-hand panel. While the D20 model concerns the migration of rings, the DSC09 and D13 models apply to the migration of clumps. The KB10 and K18 models deal with global gas transport in viscous disks and may be the most relevant for us. Note that for most of the VELA disks, irrespective of whether $Q$ is assumed to be $0.68$ or $1$, both these models overpredict the disk-averaged radial inflow levels compared to that found in the simulations. Also, all five models predict radial inflows, whereas occasionally (in about $18\%$ cases), disk-averaged outflows can be seen in the VELA simulations.
@@
Figure 6: Metallicity-Based Selection of Streams: The left-hand panel shows the 2D metallicity map, $Z_{\rm 2D}$, for VELA 7 at $z=1.5$ in a face-on projection of dimensions $4R_{\rm d}\times 4R_{\rm d}$, centered on the galactic center, obtained after mass-weighted averaging the metallicites ($Z$) of the cold gas cells along the $z$-axis between $-H_{\rm d}$ to $H_{\rm d}$. The red and cyan, dashed circles have radii equal to $R_{\rm d}$ and $2R_{\rm d}$, respectively. The map highlights an overall negative metallicity gradient from small to large $r$. The right-hand panel depicts the mass-weighted histogram of $Z$ for all cold gas cells between $r=R_{\rm d}$ to $2R_{\rm d}$ and $z=-H_{\rm d}$ to $H_{\rm d}$, normalized by the total cold gas mass inside this region. The black, dashed vertical line at $Z=0.009$ indicates the $99^{\rm th}$ percentile of this distribution, which is chosen as the threshold metallicity, $Z_{\rm cut}$, to separate streams ($Z<Z_{\rm cut}$) from non-streams ($Z>Z_{\rm cut}$). The inherent assumption made in this selection procedure is that almost all the cold gas beyond $R_{\rm d}$ is in the form of recently accreted streams.
@@
Figure 7: Streamline-Based Selection of Streams: For VELA 7 at $z=1.5$, shown are pathlines of a random selection of cold gas cells within the disk, traced over the past average disk dynamical time, $t_{\rm trace}=R_{\rm d}/\langle V_{\rm rot}\rangle$, and projected onto the $x$-$y$ plane, separated into streams (left-hand panel) and non-streams (right-hand panel). For better clarity, only part of the cold gas cells within the disk are shown. The lines are color-coded by ±$\pm d$, in units of $R_{\rm d}$, where $d$ is the distance traveled by a cold gas cell during $t_{\rm trace}$, with positive values indicating cases where the 3D radius at $t-t_{\rm trace}$ is greater than the magnitude of the current 3D radius and negative values highlighting the opposite. The red, dashed circles denote the radial extent of the disk. A threshold of $d_{\rm cut}=+0.1R_{\rm d}$ is chosen to distinguish between incoming streams and non-streams. Most of the cells classified as streams (left-hand panel) fall spatially into two distinct groups, both with large positive $d$ values, inflowing into the galaxy either from the top or bottom-right of the image. The cells classified as non-streams (right-hand panel) either have little or no displacement over $t_{\rm trace}$ ($d\approx 0$) or are outflowing (large negative $d$ values).
@@
Figure 8: Stream Maps: The left-, middle, and right-hand panels, respectively, show the face-on radial velocity (top row) and surface density (bottom row) maps of all cold gas, streams obtained with the metallicity-based selection, and streams obtained with the streamline-based selection for VELA 7 at $z=1.5$. For easy visual identification of streams, the maps in the left-hand panels are zoomed out to a region of dimensions $3R_{\rm d}\times 3R_{\rm d}$, while the actual classification of cold gas cells into streams and non-streams is only done within the disk (i.e., within a cylinder of radius $R_{\rm d}$ and height $H_{\rm d}$), as shown in the middle and right-hand panels. The red, dashed circles indicate the radial extent of the disk. From the left-hand panels, visually, there appear to be two distinct groups of incoming cold gas (blue indicates inflows in the $v_{r,\rm 2D}$ map), one entering within $R_{\rm d}$ from the top and the other from the bottom-right of the maps, which then spiral onto the galaxy counter-clockwise. The metallicity- and streamline-based selections identify somewhat different parts of these groups as streams. The regions marked with the orange ellipses (a, b, and c) are the primary sources of disagreement between the two methods of classification, which is discussed in §5.2.
@@ 4.4 Comparison with Simulations
In Fig. 5, the predictions for the radial inflow velocity from the models discussed above are compared against the results from the VELA simulations. For each snapshot in our disk galaxy sample, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ is plotted against $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ with the blue, round points. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ bins of varying widths, as indicated by the horizontal error bars, with the locations of the squares along the $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ axis being at the bin-averages. The different curves show the inflow velocity predictions from the disk-instability-driven models of radial transport, as labeled, with $Q=0.68$ (marginal instability condition for thick disks; Goldreich & Lynden-Bell, 1965) in the left-hand panel and $Q=1$ (marginal instability condition for thin disks) in the right-hand panel.
@@
As noted earlier, the D20 model specifically concerns the migration of thin rings, but the fraction of VELA disks with strong rings is small ($\lesssim 30\%$, see D20, ). The DSC09 and D13 models apply to the migration of clumps. However, we have verified that the radial mass flux in the VELA disks is dominated by the off-clump material. Therefore, the KB10 and K18 models, which deal with global gas transport in viscous disks, may be the most relevant for us. From Fig. 5, we can safely conclude that for most of the VELA disks, irrespective of whether $Q$ is assumed to be $0.68$ or $1$, both these models overpredict the magnitudes of the disk-averaged radial inflow velocities compared to that found in the simulations. We also emphasize that all five models predict radial inflows, whereas occasionally (in about $18\%$ cases), disk-averaged outflows can be seen in the VELA simulations.
@@ 5 Streams
As seen in Fig. 2, the VELA galaxies are fed by non-axisymmetric streams of cold gas, which can penetrate well within the disk. As such, much of the radial mass flux and velocity associated with the VELA disks could well be dominated by the motion of recently accreted streams rather than being generated by disk instability. To ascertain if this is indeed the case, we first discuss how streams are identified in the VELA disks (§5.1) and then quantify the radial transport of streams and off-stream cold gas, labeled as non-streams, separately. We highlight a particular example in §5.2 and discuss the overall statistics from the VELA disk galaxy sample in §5.3.

### v2

@@ 3.2 General Trends
Having discussed a particular case, we now describe some general trends. In Fig. 3, the left- and right-hand panels, respectively, show the medians (solid lines) and $16^{\rm th}-84^{\rm th}$ percentiles (envelopes) of cold gas $V_{r}$ (in units of $V_{\rm rot}$) and $F_{r}$ (in units of $\mathscr{M}/t_{\rm dyn}$) as functions of $r$ (in units of $R_{\rm d}$) obtained from our disk galaxy sample. The median $V_{r}$ and $F_{r}$ are always negative and increase in magnitude with increasing $r$, except within $\sim\!0.2\ R_{\rm d}$, where the trends are reversed. While both $F_{r}$ and $V_{r}$ are positive at the $84^{\rm th}$ percentile for all $r$, their magnitudes at the $16^{\rm th}$ percentile are much larger. For example, at $r=R_{\rm d}$, $V_{r}$ at the $16^{\rm th}$ percentile is about $0.25\ V_{\rm rot}$, while that at the $84^{\rm th}$ percentile is only $\sim\!0.04\ V_{\rm rot}$. Similarly, at the same radius, $F_{r}$ at the $16^{\rm th}$ percentile is about five times larger in magnitude than that at the $84^{\rm th}$ percentile. Therefore, in a statistical sense, the radial transport of cold gas in our disk galaxy sample is inflow-dominated at all radii, i.e., the inflowing flux is larger in magnitude than the outflowing flux. This is also shown explicitly in the right-hand panels of Figure 13.
In Fig. 4, the ratio of the disk-averaged radial velocity to the disk-averaged rotational velocity, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$, is plotted against $f$ and $z$ in the left- and right-hand panels, respectively, with the blue, round points. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $f$ and $z$ bins of varying widths (to include a sensible number of snapshots), as indicated by the horizontal error bars, with the locations of the squares along the $f$- and $z$-axes being at the bin-averages. We find that for most of the VELA disks ($82\%$), $\langle V_{r}\rangle$ is negative, again highlighting the inflow-dominated nature of cold gas radial transport in our disk galaxy sample (note that $\langle V_{\rm rot}\rangle$ is always positive). Also, on average, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ decreases (increases in magnitude) with an increase in both $f$ and $z$, indicating that galaxies with larger cold gas fractions or at higher redshifts have higher levels of disk-averaged radial inflow in general. Both trends are relatively weak, though the one with cold gas fraction is more pronounced than that with redshift, the Spearman’s rank correlation coefficients in the two cases being 0.26-0.26 and 0.13-0.13, respectively.
If the in-disk radial motions are driven by internal disk instabilities, a strong correlation is expected with gas fraction (and redshift, as gas fractions increase with increasing redshift). That we only see a weak correlation hints against this scenario, which is explored in more details in the next section. Trend in the radial motion with total cold gas mass in explored in Appendix D.
@@ 4.1 Viscous Disk Transport
Here, $\mathscr{L}$ is the rate at which turbulent kinetic energy per unit area is dissipated, $V$ is the circular velocity at radius, $r$, $\beta=d{\rm ln}V/d{\rm ln}r$, $\sigma$ is the 11D gas velocity dispersion, assumed to be isotropic, and $\sigma_{\rm sf}$ is the value of $\sigma$ that can be maintained with star formation feedback alone without any additional energy input from radial transport of gas ($\dot{M}_{\rm in}=0$). Furthermore, $\dot{M}_{\rm in}=2\pi r\Sigma V_{\rm in}$, where $\Sigma$ and $V_{\rm in}$ are the gas surface density and inflow velocity, respectively. Therefore,
@@
where $\sqrt{\sigma^{2}-\sigma^{2}_{\rm th}}$ is the 11D turbulent velocity dispersion of gas, $\sigma_{\rm th}$ being the 11D thermal velocity dispersion, typically much smaller than $\sigma$, and $t_{\rm cross}=h/\sqrt{\sigma^{2}-\sigma^{2}_{\rm th}}$. The gas scale height is estimated following Forbes et al. (2012) as
@@
where $Q$ is the two-component Toomre-Q parameter (see Romeo & Falstad, 2013), expressed as $Q=f_{g,Q}Q_{g}$, where $f_{g,Q}$ depends on the gas and stellar surface densities and velocity dispersions (see Equation 9 in K18, ), $\phi_{Q}=1+Q_{g}/Q_{*}$, and $\phi_{\rm nt}=1-\sigma_{\rm th}^{2}/\sigma^{2}$. Using Equations 9 and 11,
@@ 4.2 Ring Migration
where $\sigma$ is the average 11D gas velocity dispersion, assumed to be isotropic, the angular velocity is approximated as
@@ 4.3 Clump Migration
Figure 5: Comparison of Disk Instability Based Radial Transport Models with Simulations: $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ versus $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ is plotted for each snapshot in our disk galaxy sample, using blue circles. Red squares represent the medians within bins of $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ of varying widths, selected to ensure a reasonable number of snapshots per bin. Vertical error bars indicate the $16^{\rm th}$$84^{\rm th}$ percentile ranges, and horizontal error bars show the bin widths. The red squares are positioned at the average $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ for each bin. Overlaid curves correspond to predictions from disk-instability-based models of radial inflow, as labeled. The left-hand panel adopts $Q=0.68$ (marginal instability for thick disks; Goldreich & Lynden-Bell, 1965), and the right-hand panel uses $Q=1$ (marginal instability for thin disks). The D20 model describes inflow via ring migration, while the DSC09 and D13 models focus on clump migration. The KB10 and K18 models consider global gas transport in viscous disks and are likely the most applicable for comparison with the analysis of the VELA disks presented in this paper. It is worth noting that for most VELA disks, these two models tend to overestimate the disk-averaged radial inflow, regardless of whether $Q=0.68$ or $Q=1$ is assumed. Furthermore, all five models consistently predict radial inflows, whereas about $18\%$ of the VELA disks show disk-averaged outflows.
@@
Figure 6: Cold Gas Radial Mass Flux vs Bathtub Model Prediction: For the VELA disk galaxy sample, the cold gas radial flux mass at $R_{\rm d}$, $F_{r=R_{\rm d}}$, divided by the total cold gas mass within $R_{\rm d}$, $M_{\rm d}$, is plotted as a function of the scale factor, $a$, with the solid black line and the envelope indicating the median and $16^{\rm th}-84^{\rm th}$ percentile variation over $a$ bins of width $0.07$. The brown, dot-dashed line highlights the specific accretion rate of cold gas at the disk edge, predicted by a steady-state bathtub model, ignoring any disk-instability driven mass inflows through the disk onto the central bulge or feedback-driven outflows (see D13, ). For comparison, the cosmological specific accretion rate of baryons onto the halo, also derived in see D13, is shown with the blue, dotted line. The good agreement between the black and the brown curves highlights the importance of accreted cold gas in explaining the dynamics of VELA disks.
@@
Figure 7: Metallicity-Based Selection of Streams: The left-hand panel presents the two-dimensional metallicity distribution ($Z_{\rm 2D}$) for the VELA 7 galaxy at a redshift of $z=1.5$, shown in a face-on orientation spanning $4R_{\rm d}\times 4R_{\rm d}$ and centered on the galactic center. This map is derived by computing the mass-weighted average of the cold gas metallicities ($Z$) along the $z$-axis within the vertical range of $H_{\rm d}$ to $H_{\rm d}$. Overlaid on the map are dashed red and cyan circles representing radii of $R_{\rm d}$ and $2R_{\rm d}$, respectively. The image reveals a general decline in metallicity with increasing radial distance, indicating a negative metallicity gradient. The right-hand panel displays the mass-weighted histogram of $Z$ for cold gas cells located between $r=R_{\rm d}$ and $r=2R_{\rm d}$, and vertically between $z=-H_{\rm d}$ and $z=H_{\rm d}$, normalized by the total mass of cold gas in the specified volume. The black dashed vertical line at $Z=0.009$ marks the $99^{\rm th}$ percentile of this distribution and serves as the threshold metallicity, $Z_{\rm cut}$, used to distinguish between stream gas ($Z<Z_{\rm cut}$) and non-stream gas ($Z>Z_{\rm cut}$). This stream selection procedure is based on the assumption that the vast majority of cold gas located beyond $R_{\rm d}$ originates from recently accreted streams.
@@
Figure 8: Streamline-Based Selection of Streams: For the VELA 7 galaxy at a redshift $z=1.5$, the figure displays pathlines for a randomly selected subset of cold gas cells within the galactic disk, traced backward over an average disk dynamical timescale, defined as $t_{\rm trace}=R_{\rm d}/\langle V_{\rm rot}\rangle$, and projected onto the xxyy plane. These pathlines are shown separately for streams (left-hand panel) and non-streams (right-hand panel). To enhance visual clarity, only a portion of the cold gas cells within the disk are included. The pathlines are color-coded by $\pm d$, where $d$ represents the distance a gas cell has traveled during $t_{\rm trace}$, normalized by $R_{\rm d}$. Positive values of $d$ indicate inward motion (i.e., the 3D radius at $t-t_{\rm trace}$ is larger than at the present time), while negative values signify outward movement. Red dashed circles mark the radial boundary of the disk. A threshold value of $d_{\rm cut}=+0.1R_{\rm d}$ is used to distinguish inflowing streams from non-streaming gas. Most stream-classified cells (left-hand panel) cluster into two spatially coherent inflowing structures, both showing large positive $d$ values and entering the disk either from the top or bottom-right regions of the projection. In contrast, cells labeled as non-streams (right-hand panel) either show minimal displacement over $t_{\rm trace}$ ($d\approx 0$) or exhibit outward motion with significantly negative $d$ values.
@@
Figure 9: Stream Maps: The left-, middle, and right-hand panels respectively present the face-on maps of radial velocity (top row) and surface density (bottom row) in VELA 7 at $z=1.5$, corresponding to all cold gas, streams identified via the metallicity-based method, and streams identified via the streamline-based method. To aid visual recognition of streams, the maps in the left-hand panels are shown over an expanded area of $3R_{\rm d}\times 3R_{\rm d}$. However, the actual classification of cold gas into streams and non-streams is confined to the galactic disk, defined as a cylinder with radius $R_{\rm d}$ and height $H_{\rm d}$, as shown in the middle and right-hand panels. Red dashed circles denote the radial extent of this disk. In the left-hand panels, two distinct inflowing cold gas groups are apparent (blue in the $v_{r,\rm 2D}$ map indicates inward motion), entering within $R_{\rm d}$ either from the top or the bottom-right and subsequently spiraling counter-clockwise toward the galactic center. The metallicity- and streamline-based classifications select somewhat different portions of these inflowing groups as streams. The areas enclosed by the orange ellipses (labeled a, b, and c) correspond to the main regions of disagreement between the two classification methods; these discrepancies are discussed further in §5.3
@@ 4.4 Comparison with Simulations
In Fig. 5, the predictions for the radial inflow velocity from the models discussed above are compared against the results from the VELA simulations. For each snapshot in our disk galaxy sample, $\langle V_{r}\rangle/\langle V_{\rm rot}\rangle$ is plotted against $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ with the blue, round points. The red squares and the vertical error bars denote the medians and $16^{\rm th}-84^{\rm th}$ percentiles, respectively, in $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ bins of varying widths, as indicated by the horizontal error bars, with the locations of the squares along the $\langle V_{\rm rot}\rangle/\langle\sigma_{r}\rangle$ - axis being at the bin-averages. The different curves show the inflow velocity predictions from the disk-instability-driven models of radial transport, as labeled, with $Q=0.68$ (marginal instability condition for thick disks; Goldreich & Lynden-Bell, 1965) in the left-hand panel and $Q=1$ (marginal instability condition for thin disks) in the right-hand panel.
@@
As noted earlier, the D20 model specifically concerns the migration of thin rings, but the fraction of VELA disks with strong rings is small ($\lesssim 30\%$, see D20, ). The DSC09 and D13 models apply to the migration of clumps. However, we have verified that the radial mass flux in the VELA disks is dominated by the off-clump material. Therefore, the KB10 and K18 models, which deal with global gas transport in viscous disks, may be the most relevant for us. From Fig. 5, we can safely conclude that for most of the VELA disks, irrespective of whether $Q$ is assumed to be $0.68$ or $1$, both these models overpredict the magnitudes of the disk-averaged radial inflow velocities compared to that found in the simulations (see also Appendix C, which shows that the same is true for the locally averaged inflow at different radii as well). We also emphasize that all five models predict radial inflows, whereas occasionally (in about $18\%$ cases), disk-averaged outflows can be seen in the VELA simulations.
@@ 5 Streams
As seen in Fig. 2, the VELA galaxies are fed by non-axisymmetric streams of cold gas, which can penetrate well within the disk. As such, much of the radial mass flux and velocity associated with the VELA disks could well be dominated by the motion of recently accreted streams rather than being generated by disk instability. To ascertain if this is indeed the case, we first compare the radial mass flux at the boundary of VELA disks (in a cylindrical shell of radius $r=R_{\rm d}$, width $\Delta r=0.1R_{\rm d}$, and height, $H_{\rm d}$) to the cold gas accretion rate at the disk boundary predicted by a simple “bathtub” model (e.g., Bouché et al., 2010; Dekel & Mandelker, 2014) in steady state, where the star formation rate in the disk is balanced by the accretion rate, ignoring any mass inflows through the disk onto the central bulge or mass outflows (§5.1). We then discuss how streams are identified in the VELA disks (§5.2) and quantify the radial transport of streams and off-stream cold gas, labeled as non-streams, separately. We highlight a particular example in §5.3 and discuss the overall statistics from the VELA disk galaxy sample in §5.4.
5.1 Radial Mass Flux vs Accretion Rate from Bathtub Model
Assuming mass conservation, the total cold gas mass, $M_{\rm d}$, within a disk varies according to
$\dot{M}_{\rm d}=\dot{M}_{\rm ac}-\dot{M}_{\rm sf}\,$, (31)
where $\dot{M}_{\rm ac}$ is the accretion rate of cold gas at the disk edge and $\dot{M}_{\rm sf}$ is the star formation rate, crudely approximated as
$\dot{M}_{\rm sf}=M_{\rm d}/\tau_{\rm sf}\,$, (32)
where $\tau_{\rm sf}=t_{\rm ff}/\epsilon_{\rm sf}$. Here, $t_{\rm ff}$ is the free-fall time, and $\epsilon_{\rm sf}$ is the star formation efficiency. Note that while deriving Equation 31, we have neglected all sink terms other than star formation, such as disk-instability driven mass inflows through the disk onto the central bulge and outflows driven by stellar feedback.
If $\dot{M}_{\rm ac}$ and $\tau_{\rm sf}^{-1}$ are slowly varying with timescales that are longer than $\tau_{\rm sf}$, Equations 31 and 32 can be easily solved to give
$M_{\rm d}=\dot{M}_{\rm ac}\tau_{\rm sf}(1-e^{-t/\tau_{\rm sf}})\,,\ \ \ \dot{M}_{\rm d}=\dot{M}_{\rm ac}e^{-t/\tau_{\rm sf}}\,.$ (33)
After a transition period of order $\tau_{\rm sf}$, the above equation relaxes to a steady-state solution given by
$M_{\rm d}\simeq\dot{M}_{\rm ac}\tau_{\rm sf}\,,\ \ \ \dot{M}_{\rm d}\simeq 0\,.$ (34)
Therefore, $\dot{M}_{\rm ac}/M_{\rm d}\simeq\tau_{\rm sf}^{-1}=\epsilon_{\rm sf}t_{\rm ff}^{-1}$.
Following D13, we estimate $t_{\rm ff}$ as $t_{\rm ff}=f_{\rm sf}t_{\rm d}$, where $f_{\rm sf}\sim 0.5$, assuming that star formation occurs in clumps where the overdensity is a factor of few with respect to the average disk density and
$t_{\rm d}=R_{\rm d}/V\simeq\upsilon^{-1}\lambda R_{\rm v}/V_{\rm v}\simeq 0.14\upsilon^{-1}\lambda t$ (35)
is the disk dynamical time. Here, $\lambda$ is the halo spin parameter, relating the disk radius, $R_{\rm d}$, to the halo virial radius, $R_{\rm v}$, $\upsilon$ is the ratio between the disk rotation velocity, $V$, and the halo virial velocity, $V_{\rm v}$, and using the redshift dependent relations for $R_{\rm v}$ and $V_{\rm v}$, $R_{\rm v}/V_{\rm v}=0.14t$, where $t$ is the cosmological time. Adopting $\lambda\sim 0.05$, $\upsilon\sim 1$, and $\epsilon_{\rm sf}\sim 0.02$, we get
$\dot{M}_{\rm ac}/M_{\rm d}\simeq\frac{1}{0.175t}=\frac{a^{-3/2}}{0.175t_{1}}\,$, (36)
where we have used the relation $1/a=(t/t_{1})^{-2/3}$, with $t_{1}=17.5\ {\rm Gyr}$ (see D13, ). For comparison, the cosmological specific accretion rate of baryons at the halo boundary is given by
$\frac{\dot{M}_{\rm b}}{M_{\rm b}}\simeq sa^{-5/2}\,$, (37)
with $s=0.03\ {\rm Gyr}^{-1}$, which was also derived in D13 and shown to be in good agreement with earlier VELA-like simulations.
In Fig. 6, the cold gas radial mass flux at $R_{\rm d}$, $F_{r=R_{\rm d}}$, divided by $M_{\rm d}$ is shown as a function of the scale factor $a$ for our disk galaxy sample, with the solid black line and the envelope indicating the median and the $16^{\rm th}-84^{\rm th}$ percentile variation over $a$ bins of width $0.07$. The brown, dot-dashed line plots $\dot{M}_{\rm ac}/M_{\rm d}$ as a function of $a$ according to Equation 36. For comparison, the blue, dotted line also indicates $\dot{M}_{\rm b}/M_{\rm b}$ as a function of $a$ according to Equation 37. The good agreement between the median trend in $F_{r=R_{\rm d}}/M_{\rm d}$ and $\dot{M}_{\rm ac}/M_{\rm d}$ from the bathtub model, shows that on average, the cold gas radial mass flux at the edge of VELA disks is consistent with that expected from accretion onto the disk. In the following subsections, we show that in VELA disks, the accreted cold gas streams can penetrate well within the disk, such that the radial mass flux in the outer disks ($r\gtrsim 0.5R_{{\rm d}}$) is largely dominated by stream motions.